OpenStudy (anonymous):

in an automotive part manufacturing company, the produced parts are placed in two boxes A and B , it turns that A box contain no deffective part while 1/5 parts in B are defective, a person picks up a box , suppose A and B have equal chance of selection . what is the probability that the person picks up the defective part ?

OpenStudy (anonymous):

With these probability problems, I believe you multiply probability by probability to get the final probability. Hopefully these are not homework problems being posted just to get an answer. So how would you start?

OpenStudy (anonymous):

Probability of defective part = P(D) = P(D|A)*P(A)+P(D|B)*P(B) = 0 + 1/5*1/2 =10 %

OpenStudy (anonymous):

its not homework i just want to share something , i am studying this right now :P

OpenStudy (anonymous):

Correct. 0.5 * 0.2 = 0.1 :) Ok just making sure :)